OpenStudy (anonymous):

When Fritz drives to work his trip takes 48 minutes, but when he takes the train it takes 40 minutes. Find the distance Fritz travels to work if the train travels an average of 6 miles per hour faster then his driving.

OpenStudy (anonymous):

let fritz's speed be x mph trains speed is x+6 mph x = d / t = d / 5/4 = 5d/4 x + 6 = d / 2/3 = 3d/2 solve these two equations to find x and distance d substitute x = 5d/4 into second equation 5d/4 + 6 = 3d/2 multiply thru by 4: 5d+ 24 = 6d d = 24 distance is 24 miles

OpenStudy (anonymous):

ok alma?